# Stolz positive scalar curvature structure groups, proper actions and equivariant 2-types

### v1

@@ 5.2 Construction of a universal space
We assume as induction hypothesis that for each subgroup $H$ of $G$ for each component of $X_{k-1}^{H}$ the $\pi_{j}$-th homotopy groups vanishes for $22\leq j<k-$ 2. Note that for $k=3$ this is an empty condition and hence the induction start for $k=3$ is trivially satisfied.
Then, for each subgroup $H$ of $G$ we first attach cells $H/H\times D^{k}$ of dimension $k$ to $X_{k-1}^{H}$ in order to make $\pi_{k-1}(X^{H})$ trivial. To actually remain in the world of $G$-CW-complexes we induce these attaching constructions up and attach $G/H\times D^{k}$ to $G\cdot X_{k-1}^{H}\subset X_{k-1}$. Note that these $G$-$k$-cells affect also other fixed point sets, but there will be no change of $\pi_{j}$ for $j<k-1$, and $\pi_{k-1}$ can only get smaller, but as we make sure that it is trivial, this property will not be affected. After attaching all these $G$-cells we therefore get a new $G$-CW-complex $X_{k}$ containing $X_{k-1}$ such that $\pi_{j}(X_{k}^{H},x)=0$ for all subgroups $H$ of $G$ and for all $12\leq j\leq k-$ 1 and for all basepoints $x\in X_{k}^{H}$. The union of all $X_{k}$ (with the colimit topology) is then called $B\Pi$. It contains $X$ as a subcomplex and has the same $2$-skeleton as $X$, and satisfies that $\pi_{j}(B\Pi^{H},x)=0$ for all subgroups $H$ of $G$ and all $x\in B\Pi^{H}$.
Example 5.5.
Consider the 2-dimensional torus $\mathbb{T}^{2}=\mathbb{S}^{1}\times\mathbb{S}^{1}$ with the following CW-complex structure: the first factor is given by two vertices $v_{1}$ and $v_{2}$; two edges $e_{1}$ and $e_{2}$ both with extremes $v_{1}$ and $v_{2}$ (this gives the first factor); two loops $l_{1}$ and $l_{2}$ attached to $v_{1}$ and $v_{2}$ respectively (these give two copies of the second factor); finally two 2-cells $c_{1}$ and $c_{2}$ suitably attached, the first one to $e_{1}$, $l_{1}$ and $l_{2}$, the second one to $e_{2}$, $l_{1}$ and $l_{2}$. Let then $\mathbb{Z}_{2}$ act on $\mathbb{T}^{2}$ by flipping the first factor and then fixing $l_{1}$ and $l_{2}$. Because the fixed-point sets are the aspherical spaces $\mathbb{S}^{1}$ for $\mathbb{Z}/2$ and $\mathbb{T}^{2}$ for $\{e\}$, this it is a classifying space for the functor
$\Pi\colon\operatorname{Orb}(\mathbb{Z}_{2})\to Groupoids$
which assigns to $\mathbb{Z}_{2}/\mathbb{Z}_{2}$ the groupoid ×$\mathbb{Z}\times Pair\{v_{1},v_{2}\}$ over $\{v_{1},v_{2}\}$, to $\mathbb{Z}_{2}/\{0\}$ the group $\mathbb{Z}^{2}$ and to the unique non-identity morphism in $\operatorname{Hom}_{\operatorname{Orb}(\mathbb{Z}_{2})}$ the morphism of groupoids which sends ×$\mathbb{Z}\times Pair\{v_{1},v_{2}\}$ to the same copy of $\mathbb{Z}$ in $\mathbb{Z}^{2}$, which is induced by the inclusion of the two loops $l_{1},l_{2}$ into the torus $\mathbb{T}^{2}$.
It is a nice exercise to carry out the construction of Proposition 5.4 for this $\Pi$. One will observe that one attaches quite a few more $2$-cells and that way produces a space $X$ with non-trivial second homotopy group, so not homotopy equivalent to $\mathbb{T}^{2}$.
@@
We begin by defining the map $\varphi$ on the $G$-equivariant 00-skeleton of $Y$ by putting
@@
Now, we proceed to define $\varphi$ on the $1$-skeleton of $Y$. For a $G$ $1$-cell $c$ of the form $G/H\times[0,1]$ (with isotropy $H$) pick the ordinary $1$-cell $c_{H}:=H/H\times D^{1}$ contained in $Y^{H}$. Then $c_{H}$ defines (if we choose an orientation of it) an element $\gamma\in\Pi_{1}(Y^{H})_{|Y^{H}_{(0)}}$. Define now $\phi\colon c_{H}\to(B\Pi^{H})^{(1)}$ such that this map represents the element $\Phi(H)(\gamma)$ in $\Pi_{1}(B\Pi^{H})|_{(B\Pi^{(0)})^{H}}$. Since $H$ acts trivially on the target, this extends uniquely to a $G$-map on the $G$-cell $c$ with values in the $1$-skeleton of $B\Pi$. Of course, we could have picked a different base cell in the $G$-cell $c$. But because $\Phi$ is a natural transformation, the resulting map is independent of this —up to the choice of the representative in the homotopy class of $\Phi(H)(\gamma)$ which had already to be made anyway. This defines the map $\phi$ on the $1$-skeleton of $Y$.
To extend $\phi$ to $Y^{(2)}$, let $c$ be a $G$-$2$-cell of the form $G/H\times D^{2}$. Pick the corresponding ordinary $2$-cell $H/H\times D^{2}$ (with isotropy $H$) which is contained in $Y^{H}$. Its attaching map $\psi\colon S^{1}\to Y^{H}$ is obviously contractible in $Y^{H}$ and hence trivial (i.e. a unit) in the fundamental groupoid $\Pi_{1}(B\Pi^{H})|_{(B\Pi^{(0)})^{H}}$ (we conjugate with a path in $Y^{H}$ to the 00-skeleton to make $\psi$ represent an element of $\Pi_{1}(Y^{H})_{B\Pi^{H}_{(0)}}$, the triviality does not depend on the choice of such a path). Consequently, $\Phi(H)$ being a map of groupoids, also the image under $\Phi(H)$ of this element of the fundamental groupoid is trivial (a unit). By construction of $\phi$ on the $1$-skeleton, this image element is represented by $\phi\circ\psi$ (up to the chosen conjugation with a path to the 00-skeleton), which is hence contractible in $B\Pi^{H}$. Extend $\phi$ over $H/H\times D^{2}$ using this contraction and then extend it $G$-equivariantly to $c=G/H\times D^{2}$.
@@
Observe that when $G=\{e\}$ is the trivial group and $\Pi(\{e\}$ is a discrete group $\Gamma$ (a groupoid with only one object) then $B\Pi$ is a standard $CW$-complex with a single 00-cell. It follows that the space obtained in this way is an Eilenberg-Mac Lane space $B\Gamma=K(\Gamma,1)$.
@@
Recall the standard two fact:
@@
[…]
@@
Let $\Pi\colon\operatorname{Orb}(G)\to Groupoids$ be a functor (though of as an abstract fundamental groupoid functor). Let $X$ be a $G$-CW complex and $\Phi\colon\Pi_{1}(X;G)\to\Pi$ be an equivalence as in Definition 5.8.
Then for the classifying $G$-map $\varphi\colon X\to B\Pi$ as in Proposition 5.6 which induces $\Phi$ on the level of fundamental groupoid functors, we have that
@@
By Lemma 5.11, the map $f\colon X\to Y$ induces isomorphisms of $\pi_{0}$ and $\pi_{1}$ for each fixed point set $X^{H}$ and each basepoint. Moreover, as all higher homotopy groups of fixed point sets of $B\Pi$ are trivial, this map indeed is $2$-connected in the sense of 2.3. Therefore, the assertion is exactly the one of Theorem 4.1. ∎
@@
Let $f\colon X\to Y$ be a $G$-map between $G$-CW-complexes such that for each subgroup $H$ of $G$ the induced map $f^{H}\colon X^{H}\to Y^{H}$ induces an isomorphism $\pi_{j}(f^{H})\colon\pi_{j}(X^{H},x)\to\pi_{j}(Y^{H},f(x))$ for all $x\in X^{H}$ and $j=0,1$. Then

### v2

@@ 5.2 Construction of a universal space
We assume as induction hypothesis that for each subgroup $H$ of $G$ for each component of $X_{k-1}^{H}$ the $\pi_{j}$-th homotopy groups vanishes for $2\leq j<k-2$. Note that for $k=3$ this is an empty condition and hence the induction start for $k=3$ is trivially satisfied.
Then, for each subgroup $H$ of $G$ we first attach cells $H/H\times D^{k}$ of dimension $k$ to $X_{k-1}^{H}$ in order to make $\pi_{k-1}(X^{H})$ trivial. To actually remain in the world of $G$-CW-complexes we induce these attaching constructions up and attach $G/H\times D^{k}$ to $G\cdot X_{k-1}^{H}\subset X_{k-1}$. Note that these $G$-$k$-cells affect also other fixed point sets, but there will be no change of $\pi_{j}$ for j<k1j<k-1, and $\pi_{k-1}$ can only get smaller, but as we make sure that it is trivial, this property will not be affected. After attaching all these $G$-cells we therefore get a new $G$-CW-complex $X_{k}$ containing $X_{k-1}$ such that $\pi_{j}(X_{k}^{H},x)=0$ for all subgroups $H$ of $G$ and for all $2\leq j\leq k-1$ and for all basepoints $x\in X_{k}^{H}$.
Definition 5.5.
The union of all $X_{k}$ (with the colimit topology) is then called $B\Pi$.
It has the following characteristic property:
Lemma 5.6.
The $G$-space $B\Pi$ just constructed contains $X$ as a subcomplex and has the same $2$-skeleton as $X$, and satisfies that $\pi_{j}(B\Pi^{H},x)=0$ for all $j\in\mathbb{N}$ and for all subgroups $H$ of $G$ and all $x\in B\Pi^{H}$.
@@
We begin by defining the map $\varphi$ on the $G$-equivariant 0-skeleton of $Y$ by putting
@@
Now, we proceed to define $\varphi$ on the $1$-skeleton of $Y$. For a $G$ 11-cell $c$ of the form $G/H\times[0,1]$ (with isotropy $H$) pick the ordinary $1$-cell $c_{H}:=H/H\times D^{1}$ contained in $Y^{H}$. Then $c_{H}$ defines (if we choose an orientation of it) an element $\gamma\in\Pi_{1}(Y^{H})_{|Y^{H}_{(0)}}$. Define now $\phi\colon c_{H}\to(B\Pi^{H})^{(1)}$ such that this map represents the element $\Phi(H)(\gamma)$ in $\Pi_{1}(B\Pi^{H})|_{(B\Pi^{(0)})^{H}}$. Since $H$ acts trivially on the target, this extends uniquely to a $G$-map on the $G$-cell $c$ with values in the $1$-skeleton of $B\Pi$. Of course, we could have picked a different base cell in the $G$-cell $c$. But because $\Phi$ is a natural transformation, the resulting map is independent of this —up to the choice of the representative in the homotopy class of $\Phi(H)(\gamma)$ which had already to be made anyway. This defines the map $\phi$ on the $1$-skeleton of $Y$.
To extend $\phi$ to $Y^{(2)}$, let $c$ be a GG-22-cell of the form $G/H\times D^{2}$. Pick the corresponding ordinary $2$-cell $H/H\times D^{2}$ (with isotropy $H$) which is contained in $Y^{H}$. Its attaching map $\psi\colon S^{1}\to Y^{H}$ is obviously contractible in $Y^{H}$ and hence trivial (i.e. a unit) in the fundamental groupoid $\Pi_{1}(B\Pi^{H})|_{(B\Pi^{(0)})^{H}}$ (we conjugate with a path in $Y^{H}$ to the 0-skeleton to make $\psi$ represent an element of $\Pi_{1}(Y^{H})_{B\Pi^{H}_{(0)}}$, the triviality does not depend on the choice of such a path). Consequently, $\Phi(H)$ being a map of groupoids, also the image under $\Phi(H)$ of this element of the fundamental groupoid is trivial (a unit). By construction of $\phi$ on the $1$-skeleton, this image element is represented by $\phi\circ\psi$ (up to the chosen conjugation with a path to the 0-skeleton), which is hence contractible in $B\Pi^{H}$. Extend $\phi$ over $H/H\times D^{2}$ using this contraction and then extend it $G$-equivariantly to $c=G/H\times D^{2}$.
@@
Example 5.8.
Consider the 2-dimensional torus $\mathbb{T}^{2}=\mathbb{S}^{1}\times\mathbb{S}^{1}$ with the following CW-complex structure: the first factor is given by two vertices $v_{1}$ and $v_{2}$; two edges $e_{1}$ and $e_{2}$ both with extremes $v_{1}$ and $v_{2}$ (this gives the first factor); two loops $l_{1}$ and $l_{2}$ attached to $v_{1}$ and $v_{2}$ respectively (these give two copies of the second factor); finally two 2-cells $c_{1}$ and $c_{2}$ suitably attached, the first one to $e_{1}$, $l_{1}$, $e_{1}^{-1}$, and $l_{2}$, the second one to $e_{2}$, $l_{1}$, $e_{2}^{-1}$, and $l_{2}$. Let then $\mathbb{Z}_{2}$ act on $\mathbb{T}^{2}$ by flipping the first factor (in particular, swapping $e_{1}$ and $e_{2}$) and also $c_{1}$ and $c_{2}$, and then fixing $l_{1}$ and $l_{2}$.
Figure 1: Torus with action of $\mathbb{Z}_{2}$
In Figure 1 we represent the torus as a square where we identify opposite sides. Because the fixed-point sets are the aspherical spaces $\mathbb{S}^{1}$ for $\mathbb{Z}_{2}$ and $\mathbb{T}^{2}$ for $\{e\}$, this is a classifying space for the functor $\Pi_{1}(T^{2};\mathbb{Z}_{2})$, isomorphically given as
$\Pi\colon\operatorname{Orb}(\mathbb{Z}_{2})\to Groupoids$
which is defined as follows. We have to specify which groupoid it assigns to each of the two possible $\mathbb{Z}_{2}$-orbits, namely to the trivial orbit $\mathbb{Z}_{2}/\mathbb{Z}_{2}$ and the free orbit $\mathbb{Z}_{2}/\{0\}$. Moreover, in the category $\operatorname{Orb}(\mathbb{Z}_{2})$ there are only two non-identity morphisms, the collapse map $\mathbb{Z}_{2}/\{0\}\to\mathbb{Z}_{2}/\mathbb{Z}_{2}$ and the non-identity bijection $\tau\colon\mathbb{Z}_{2}/\{0\}\to\mathbb{Z}_{2}/\{0\}$ and we have to specify the associated morphisms of groupoids.
Concretely, we define $\Pi$ as follows.
  • We assign to $\mathbb{Z}_{2}$ the groupoid $\mathbb{Z}\times\{v_{1},v_{2}\}$ over $\{v_{1},v_{2}\}$, where $\{S\}$ denotes the trivial groupoid of the set $S$ with only identity morphisms.
  • We assign to $\{0\}\subset\mathbb{Z}_{2}$ the group $\mathbb{Z}^{2}$.
  • We assign to the collapse map $\mathbb{Z}_{2}/\{0\}\to\mathbb{Z}_{2}/\mathbb{Z}_{2}$ the morphism of groupoids which first projects $\mathbb{Z}\times\{v_{1},v_{2}\}$ to the factor $\mathbb{Z}$ and then injects $\mathbb{Z}$ into $\mathbb{Z}^{2}$ as the subgroup $\{0\}\times\mathbb{Z}$. Note that, if we identify $\mathbb{Z}^{2}$ with the fundamental group of the torus $T^{2}$ we constructed above, we think of this as the subgroup generated by each of the two loops $l_{1}$, or $l_{2}$ (the two loops are homotopic).
  • We assign to the non-identity map $\tau\colon\mathbb{Z}_{2}/\{0\}\to\mathbb{Z}_{2}/\{0\}$ the automorphism of $\mathbb{Z}^{2}$ which sends the generator of the first factor to its inverse and the second generator to itself.
Note that indeed this is precisely the fundamental groupoid of the $\mathbb{Z}_{2}$-space $T^{2}$ with the chosen CW-structure, and by the contractibility of $S^{1}$ and $T^{2}$ it is a model for $B\Pi_{1}(T^{2};\mathbb{Z}_{2})$.
It is a nice exercise to carry out the construction of the space $X$ of Proposition 5.4 for this $\Pi$. One observes that one attaches quite a few $2$-cells and $3$-cells and that way produces a space $X$ not homotopy equivalent to $\mathbb{T}^{2}$. Specifically, the 0-skeleton of $X$ is the 0-skeleton of $T^{2}$. To obtain the $2$-skeleton, we use the classifying spaces $T^{2}$ of $\Pi(\{0\})=\mathbb{Z}^{2}$ and $S^{1}\times\{v_{1},v_{2}\}$ of $\Pi(\mathbb{Z}_{2})=\mathbb{Z}\times\{v_{1},v_{2}\}$.
By construction of $X$, we have to take the disjoint union of the $2$-skeleta of $\mathbb{Z}_{2}/\{0\}\times T^{2}$ and of $\mathbb{Z}_{2}/\mathbb{Z}_{2}\times S^{1}\times\{v_{1},v_{2}\}$ which is the disjoint union of two copies of $T^{2}$ and of $S^{1}$. To this, we have to glue $4$ cylinders for the $4$ morphisms in $\operatorname{Orb}(\mathbb{Z}_{2})$
$\begin{split}&\mathbb{Z}_{2}/\{0\}\times[0,1]\times T^{2},\qquad\mathbb{Z}_{2}/\{0\}\times[0,1]\times T^{2},\\ &\mathbb{Z}_{2}/\{0\}\times[0,1]\times S^{1}\times\{v_{1},v_{2}\},\qquad[0,1]\times S^{1}\times\{v_{1},v_{2}\}.\end{split}$
Gluing in the first cylinder, corresponding to $\operatorname{id}_{\mathbb{Z}_{2}/\{0\}}$, produces $\mathbb{Z}_{2}/\{0\}\times T^{3}$. The second cylinder produces another copy of $T^{3}$, glued with the previous two along embedded copies of $T^{2}$. The third cylinder, corresponding to the collapse map, homotopically and $\mathbb{Z}_{2}$-equivariantly glues the $2$ copies of $S^{1}$ into this space (without changing the homotopy type). The forth cylinder, corresponding to the identity of $\mathbb{Z}_{2}/\mathbb{Z}_{2}$, glues in two more copies of $T^{2}$ into the space obtained so far along a homotopically non-trivial circle in each. From the cellular chain complex we can read off that $H^{3}(X;\mathbb{Z})\cong\mathbb{Z}^{3}$, generated by the fundamental classes of the three copies of $T^{3}$ the first two cylinders produced. In particular, this space is not homotopy equivalent to $T^{2}$.
@@
Observe that when $G=\{e\}$ is the trivial group and $\Pi(\{e\})$ is a discrete group $\Gamma$ (a groupoid with only one object) then $B\Pi$ is a standard $CW$-complex with a single 0-cell. It follows that the space obtained in this way is an Eilenberg-Mac Lane space $B\Gamma=K(\Gamma,1)$.
@@
Recall the standard two facts:
@@
${\displaystyle\bigsqcup_{[x]\in\pi_{0}(\mathcal{G})}\mathcal{G}^{x}_{x}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}\scriptstyle{\simeq}$${\displaystyle\bigsqcup_{[x]\in\pi_{0}(\mathcal{H})}\mathcal{H}^{T(x)}_{T(x)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\mathcal{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{T}\textstyle{\mathcal{H}}$
@@
Let $\Pi\colon\operatorname{Orb}(G)\to Groupoids$ be a functor (thought of as an abstract fundamental groupoid functor). Let $X$ be a $G$-CW complex and $\Phi\colon\Pi_{1}(X;G)\to\Pi$ be an equivalence as in Definition 5.10.
Then for the classifying $G$-map $\varphi\colon X\to B\Pi$ as in Proposition 5.7 which induces $\Phi$ on the level of fundamental groupoid functors, we have that for $n\geq 6$
@@
By Lemma 5.13, the map $f\colon X\to Y$ induces isomorphisms of $\pi_{0}$ and $\pi_{1}$ for each fixed point set $X^{H}$ and each basepoint. Moreover, as all higher homotopy groups of fixed point sets of $B\Pi$ are trivial, this map indeed is $2$-connected in the sense of 2.6. Therefore, the assertion is exactly the one of Theorem 4.1. ∎
@@
Let $f\colon X\to Y$ be a $G$-map between $G$-CW-complexes such that for each subgroup $H$ of $G$ the induced map $f^{H}\colon X^{H}\to Y^{H}$ induces an isomorphism $\pi_{j}(f^{H})\colon\pi_{j}(X^{H},x)\to\pi_{j}(Y^{H},f(x))$ for all $x\in X^{H}$ and $j=0,1$. Then for $n\geq 6$ the induced map
@@
Remark 5.16.
Note that Proposition 5.14 and Corollary 5.15 of course also hold if $G$ is trivial and just state that $R_{n}^{spin}(X)$ depends only on the fundamental group of the CW-complex $X$.
Acknowledgements
The authors thank the referees for the careful reading of the paper, improving the presentation and helping to avoid inaccuracies.